If 4 cards are dealt from a standard deckof 52 playing cards, what is the probability that 

3 will be diamonds?

Using Hypergeometric Distribution


1
Expert's answer
2022-04-11T15:49:17-0400

Suppose that "X" is a random variable that has a hypergeometric distribution with parameters "N", "K". "N" is a size of population, "K" is a number of objects of a certain type. "n" is a number of draws and "k" is a number of successes(the objects of a certain type). We have: "P(X=k)=C_K^k\frac{C_{N-K}^{n-k}}{C_N^n}," where "C_i^j" denotes a binomial coefficient. In our case we have: "N=52", "K=13", "n=4", "k=3." Thus, "P(X=k)=C_{13}^3\frac{C_{39}^{1}}{C_{52}^4}=\frac{13!}{3!10!}\frac{39}{\frac{52!}{4!48!}}\approx0.0412"